OpenStudy (anonymous):

7 log base 9 (x+8)=7

OpenStudy (anonymous):

divide by 7 to get log9 (x+8) = 1 change to exponential form 9^1 = x+8 now solve for x....

OpenStudy (anonymous):

When b raised to the power of y is equal x: b y = x Then base b logarithm of x is equal to y: logb(x) = y For example when: 24 = 16 Then log2(16) = 4 Logarithm as inverse function of exponential function The logarithmic function, y = logb(x) is the inverse function of the exponential function, x = by So if we calculate the exponential function of the logarithm of x (x>0), f (f -1(x)) = blogb(x) = x Or if we calculate the logarithm of the exponential function of x, f -1(f (x)) = logb(bx) = x Natural logarithm (ln) Natural logarithm is a logarithm to the base e: ln(x) = loge(x) When e constant is the number:

OpenStudy (anonymous):

how did you type all that that fast?

OpenStudy (anonymous):

my typing is really very fast

OpenStudy (anonymous):

i see....

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw

OpenStudy (lgbasallote):

dpalnc is funny =))))